If we use a stiffer spring without changing the mass how will the period of oscillation change?

Grade:12th pass

1 Answers

Nishant Vora IIT Patna
7 years ago
Time period ofoscillation in spring mass system isT = 2\pi \sqrt{\frac{m}{k}}

hence if we change the stiffness k then T will change as 1/ root(k)